# Non-Hermitian ultra-strong bosonic condensation through interaction-induced caging

### v1

@@ Abstract
The single-boson band structure, shown in Fig. S1(a), contains two symmetrically gapped bulk bands ($E_{1}$ and $E_{2}$) separated by in-gap topological zero modes ($E_{0}$). The corresponding eigenstates demonstrate distinct localization patterns, with extended states for bulk bands [Fig. S1(b)] and localized states for zero modes [Fig. S1(c)].
@@
Figure S1: Single-boson spectrum of the SSH model. (a) The energy spectrum shows two bulk bands ($E_{1}$ and $E_{2}$) separated by topological zero modes ($E_{0}$). (b) Localization pattern $|\psi|^{2}$ of extended states in the bulk bands. (c) Localization of the topological zero mode, showing edge confinement.
Notably, the central band [band 3 in Fig. S2] emerges from the hybridization of two distinct configurations: one where the bosons are distributed between the lower and upper bulk bands ($E_{1}\oplus E_{2}$), and another where both bosons occupy the zero mode ($E_{0}\oplus E_{0}$). This hybridization explains why we observe five bands rather than six possible energy combinations, as the energetic proximity of these configurations leads to their mixing into a single hybridized band.
The localization patterns shown in Fig. S2(b) and (c) reflect this hybridization, exhibiting characteristics that can be understood as superpositions of the single-particle localization patterns.
S2 Possible experimental realization
In the main text, we have proposed a bosonic interacting model in a non-Hermitian SSH lattice, as shown in Eq. S1. The Hamiltonian is
@@
$+\sum_{x=1}^{L-1}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2},$
Various physical platforms have experimentally realized non-Hermitian systems, including mature mechanical platforms [100, 101, 102, 103], photonic platforms [104, 45, 105, 106, 107], electrical circuit platforms [108, 109, 61, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120], and quantum circuit platforms [13, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130]. In the meantime, multiboson physics can be explored using cold atoms in optical lattices [15, 16, 17, 18, 19], photonic waveguide arrays [131, 132, 133, 134, 135, 136, 137, 138], and superconducting circuits [25, 26, 27, 28].
Below, we propose a potential experimental realization of the bosonic interacting model, as presented in Eq. S1, using resonator arrays in real space.
@@
Arrays of micro-resonators have been extensively utilized to implement one-dimensional (1D) and two-dimensional (2D) photonic lattices, particularly in the investigation of non-Hermitian photonics over the past decade [139, 140, 141, 142, 143, 144]. These designs predominantly employ ring resonators, which can achieve exceptionally high Q factors [144].
@@
It is noted that in Ref. [105], the authors demonstrate a feasible design of a one-dimensional non-Hermitian Su-Schrieffer-Heeger model based on photonic coupled resonant arrays in Fig. S3. The non-Hermitian asymmetric coupling can be realized by the judicious design of optical gain and loss elements into unidirectional coupling link rings, as shown in Fig. S3(b). The corresponding Hamiltonian is [105]
$H(k)=\left(\begin{array}[]{cc}0&t_{1}-\frac{\gamma}{2}+t_{2}e^{ik}\\ t_{1}+\frac{\gamma}{2}+t_{2}e^{-ik}&0\end{array}\right).$ (S2)
Here, in their model, the hopping amplitudes $t_{1}\pm\gamma$ and $t_{2}$ correspond to intracell hoppings $t_{L},t_{R}$ and intercell hopping (where we have set to be 1 for simplicity) in our model in Eq. S1, respectively. The non-Hermitian skin effect is observed in their model, which is consistent with our results.
Figure S2: Two-boson spectrum in the non-interacting limit ($U=0$). (a) Energy spectrum showing five bands arising from combinations of single-particle energies. Band 1 results from hybridization of $E_{1}\oplus E_{1}$. Band 2 results from hybridization of $E_{1}(E_{0})\oplus E_{0}(E_{1})$. Band 3 results from hybridization of $E_{1}\oplus E_{2}$ and $E_{0}\oplus E_{0}$ configurations. Band 4 results from hybridization of $E_{2}(E_{0})\oplus E_{0}(E_{2})$. Band 5 results from hybridization of $E_{2}\oplus E_{2}$. (b) Localization pattern of a zero-corner state in band 3. (c) Localization of mixed bulk-topological states in bands 2 and 4.
Figure S3: Schematic of the photonic coupled resonator arrays for realizing the non-Hermitian SSH model. The judicious design of optical gain and loss elements into unidirectional coupling link rings realise the non-Hermitian asymmetric coupling coupling coupling. This figure is adapted from Ref. [105].
Based on the design in Ref. [105], we only need the nonlinearity to be implemented in the system. The Kerr effect in the micro-resonators can introduce the nonlinearity. The Kerr effect, i.e., the third-order nonlinearity $\chi^{(3)}$, is a nonlinear optical effect that occurs when the refractive index of a material changes with the intensity of light.
@@
The third-order nonlinearity can be achieved through natural materials (although the $\chi^{(3)}$ value is relatively small) or by employing cavity QED based on electromagnetically induced transparency (EIT) to achieve an equivalent $\chi^{(3)}$. The latter involves introducing atoms [79, 145, 146, 147] or quantum dots into a cavity [49, 148, 149, 150, 151], and through appropriate external field control, inducing a target optical field with strong effective nonlinearity. Additionally, the versatility of synthetic dimensions provides opportunities to implement the nonlinear term $\frac{U}{2}n_{x_{0}}^{2}$ in Eq. S1 demonstrated in Eq. (14) in Ref. [78].

### v2

@@ Abstract
The single-boson band structure, shown in Fig. S1(a), contains two symmetrically gapped bulk bands ($E_{1}$ and $E_{2}$) separated by in-gap topological zero modes ($E_{0}$) at zero energy. The corresponding eigenstates demonstrate distinct localization profiles, with boundary skin-localized states for bulk bands [Fig. S1(b)] as well as topological zero modes [Fig. S1(c)].
Figure S1: Single-boson spectrum of the non-Hermitian SSH model [Eq. S1 with $U=0$] under open boundary conditions with $t_{L}=1.6$, $t_{R}=0.4$ and $t_{0}=3$. (a) The energy spectrum shows two bulk bands ($E_{1}$ and $E_{2}$) separated by topological zero modes ($E_{0}$). (b) Localization profile $|\psi|^{2}$ of left boundary-localized states in the bulk bands. (c) Localization of the topological zero mode, also confined at the left boundary.
@@
Figure S2: Two-boson spectrum in the non-interacting limit ($U=0$). (a) Energy spectrum showing five bands arising from combinations of single-particle energies. Band 1 results from hybridization of $E_{1}\oplus E_{1}$. Band 2 results from hybridization of $E_{1}(E_{0})\oplus E_{0}(E_{1})$. Band 3 results from hybridization of $E_{1}\oplus E_{2}$ and $E_{0}\oplus E_{0}$ configurations. Band 4 results from hybridization of $E_{2}(E_{0})\oplus E_{0}(E_{2})$. Band 5 results from hybridization of $E_{2}\oplus E_{2}$. (b) Localization pattern of a zero-corner state in band 3. (c) Localization of mixed bulk-topological states in bands 2 and 4, where either $x$ or $x^{\prime}$ (both not both) are boundary-localized.
Notably, the central band [band 3 in Fig. S2] emerges from the merging of two distinct contributions: one where the bosons are distributed between the lower and upper bulk bands ($E_{1}\oplus E_{2}$), and another (much rarer) instance where both bosons occupy the zero mode ($E_{0}\oplus E_{0}$). This explains why we observe five bands rather than six possible energy band combinations, as the energetic proximity of these configurations leads to their mixing into a single band.
The localization profiles shown in Fig. S2(b) and (c) reflect this hybridization, exhibiting characteristics that can be understood as superpositions of the single-particle localization profiles.
S2 Possible physical realizations
In the main text, we have proposed a bosonic interacting model in a non-Hermitian SSH lattice, as shown in Eq.1, also reproduced here as Eq. S1:
@@
$+\sum_{x=1}^{L-1}t_{0}\left(b_{2x}^{\dagger}b_{2x+1}+b_{2x+1}^{\dagger}b_{2x}\right)+\frac{U}{2}n_{x_{0}}^{2}.$
@@
Below, we discuss how the bosonic interacting model of Eq. S1 may be potentially realized using photonic resonator arrays. Arrays of micro-resonators have been extensively utilized to implement one-dimensional (1D) and two-dimensional (2D) photonic lattices, particularly in the investigation of non-Hermitian photonics over the past decade [136, 137, 138, 139, 140, 141]. These designs predominantly employ ring resonators, which can achieve exceptionally high Q factors [141].
@@
In Ref. [142], the authors proposed how a non-Hermitian SSH model can be realized using photonic coupled resonant arrays. The non-Hermitian asymmetric coupling can be realized through judicious incorporation of optical gain and loss elements into unidirectional coupling link rings.
@@
To implement the nonlinear density term, we can employ the Kerr effect which introduces third-order nonlinearity $\chi^{(3)}$. This nonlinear optical effect causes the refractive index of a medium to change with the intensity of light. The third-order nonlinearity can be achieved through natural materials [143, 144, 145, 146] (although the $\chi^{(3)}$ value is relatively small) or by employing cavity QED based on electromagnetically induced transparency (EIT) to achieve an equivalent $\chi^{(3)}$. The latter involves introducing atoms [77, 147, 148, 149] or quantum dots into a cavity [34, 150, 151, 152, 153] and, through appropriate external field control, inducing a target optical field with strong effective nonlinearity.
S3 Bosonic condensation in the Hatano-Nelson model
In the main text, we have discussed the emergence of ultra-strong bosonic condensation in the non-Hermitian SSH model. Here, we show that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model but can also be observed in the Hatano-Nelson model. The Hatano-Nelson model is known to exhibit a non-Hermitian skin effect, where the bulk states are localized at the edges of the system [154]. After including the bosonic interactions, the interacting Hatano-Nelson Hamiltonian is given by
$H_{\text{HN}}=\sum_{x=1}^{L-1}t_{L}b_{x}^{\dagger}b_{x+1}+t_{R}b_{x+1}^{\dagger}b_{x}+\frac{U}{2}n_{x_{0}}^{2},$ (S2)
where $t_{L}$ and $t_{R}$ are the left and right hopping amplitudes, respectively. And $U$ is the on-site interaction strength. We will use the same measurement as in the main text—specifically, the smoothed two-boson correlation probability, $\bar{P}_{xx^{\prime}}(t)$, from Eq. 6 in the main text, which represents the trends of the probability of observing bosons at sites $x$ and $x^{\prime}$ simultaneously at time $t$—to demonstrate ultra-strong bosonic condensation in the Hatano-Nelson model.
In Fig. S3, we show the smoothed two-boson correlation probability $\bar{P}_{11}(t)$ and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the Hatano-Nelson model. We consider two different initial states similar as in the main text: one where both bosons are initially inside the cage $(1,1)$ and another where both bosons are initially outside the cage $(3,3)$. In both cases, we found the predominantly red curves (non-Hermitian, Interacting) exhibit the same ultra-strong bosonic condensation and caging behavior as in the Hatano-Nelson model. This demonstrates that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model but can be observed in other non-Hermitian models as well.
Figure S3: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the Hatano-Nelson model in Eq. S2.
S4 Bosonic condensation in the non-Hermitian SSH model with trivial topology
In the main text, we discussed the emergence of ultra-strong bosonic condensation in the non-Hermitian SSH model with nontrivial topology, where $t_{L}\times t_{R}=0.4\times 1.6<t_{0}^{2}=3^{2}$. Here, we demonstrate that both the ultra-condensation phenomenon and the caging mechanism can also be observed in the non-Hermitian SSH model with trivial topology, where $t^{\prime}_{L}\times t^{\prime}_{R}>t_{0}^{\prime 2}$.
For this section, we use the parameters $t_{0}^{\prime}=1$, $t^{\prime}_{L}=4.8$, and $t^{\prime}_{R}=1.2$. These values are chosen such that $t_{0}(t_{L}+t_{R})=t_{0}^{\prime}(t^{\prime}_{L}+t^{\prime}_{R})$ and $r=t^{\prime}_{L}/t^{\prime}_{R}=4$, ensuring similar scales and hopping asymmetry for both the nontrivial and trivial cases. Here, $t^{\prime}_{L}\times t^{\prime}_{R}=4.8\times 1.2>t_{0}^{\prime 2}=1$, corresponding to the trivial case. The same measurement as in the main text—specifically, the smoothed two-boson correlation probability, $\bar{P}_{xx^{\prime}}(t)$, from Eq. 6 in the main text—is used to demonstrate ultra-strong bosonic condensation in the non-Hermitian SSH model with trivial topology. And $\bar{P}_{\text{cage}}(t)$, the smoothed two-boson probability within the $x_{0}\times x_{0}$ cage, is used to illustrate the caging mechanism.
Figure S4: (a) and (b) show the smoothed two-boson correlation probability $P_{11}(t)$ in Eq. 6 in the main text and the smoothed two-boson probability $\bar{P}_{\text{cage}}(t)$ within the $x_{0}\times x_{0}$ cage for the initial state $(1,1)$ [inside the cage] at interaction position $x_{0}=2$. (c) and (d) show the same quantities for the initial state $(3,3)$ [outside the cage] and $x_{0}=2$. $r=t_{L}/t_{R}=4,t_{L}=1/t_{R}=2,U=6,L=10$ for the non-Hermitian SSH model in Eq. 1 in the main text with $t_{0}=t_{0}^{\prime}=1$, $t_{L}=t^{\prime}_{L}=4.8$, and $t_{R}=t^{\prime}_{R}=1.2$.
The ultra-strong bosonic condensation and caging behavior are illustrated in Fig. S4. We also consider two different initial states: one where both bosons are initially inside the cage $(1,1)$ and another where both bosons are initially outside the cage $(3,3)$. In both scenarios, we found that the predominantly red curves (non-Hermitian, interacting) exhibit the same ultra-strong bosonic condensation and caging behavior as observed in the non-Hermitian SSH model with nontrivial topology. This demonstrates that the ultra-condensation phenomenon and the caging mechanism are not exclusive to the non-Hermitian SSH model with nontrivial topology; rather, they can also be observed in the non-Hermitian SSH model with trivial topology.
The key difference between the trivial and nontrivial models, as discussed in the main text, lies in the absence of a topological origin in the trivial case. In this case, the topological bands, specifically bands 2 and 4, are no longer present, as illustrated in Fig. S5(b). Instead, the bulk states are localized at the edges, which is also shown in Fig. S5(b). The ultra-strong condensation described in the main text, which has a topological origin in the nontrivial case, is absent in the trivial case. Furthermore, the ultra-strong condensation shown in Fig. S5(a) does not correspond to the boundary localization of the 2-boson spectral bands in the trivial model.
Figure S5: Extent of two-boson condensation due to a boundary density interaction at $x_{0}=1$ [Eq. 1 in the main text], and its correspondence with the boundary localization of 2-boson spectral bands. (a) Time-averaged two-boson condensation probability $\bar{P}_{11}$ [Eq. 6 in the main text] at site $x=1$, in the parameter space of non-Hermitian hopping asymmetry $r$ and density interaction strength $U$. (b) Corresponding two-boson spectra at $r=4$, which features 3 bands only. The topological bands are absent, and the bulk states are localized at the edges. The interesting topological origin descibed in the main text of the ultra-strong condensation in the nontrivial case is absent in the trivial case. Parameters are $t_{0}^{\prime}=1$, $t^{\prime}_{L}=4.8$, and $t^{\prime}_{R}=1.2$.